The generic metal hydroxide M(OH)2 has Ksp = 5.65×10−12. (NOTE: In this particular problem, because of the magnitude of the Ksp and the stoichiometry of the compound, the contribution of OH− from water can be ignored. However, this may not always be the case.) Part A What is the solubility of M(OH)2 in pure water? Express your answer with the appropriate units. Part B What is the solubility of M(OH)2 in a 0.202 M solution of M(NO3)2? Express your answer with the appropriate units.
Answer #1

PART(A)

M(OH)2 (s) \rightleftharpoons M2+ (aq.) + 2 OH- (aq.)

S                         S                  2S

Ksp = [M2+][OH-]2

Ksp = (S)(2S)2

5.65 * 10-12 = 4S3

S = Solubility of M(OH)2 = 1.12 * 10-4 M

PART (B)

M(OH)2 (s) \rightleftharpoons M2+ (aq.) + 2 OH- (aq.)

M(NO3)2 (aq.) ------------> M2+ (aq.) + 2 NO3- (aq.)

0.202                                 0.202            2(0.202)

Ksp = [M2+][OH-]2

5.65 * 10-12 = (0.202)[OH-]2

[OH-]2 = 28.0 * 10-12

[OH-] = 5.29 * 10-6 M

So,

Solubility of M(OH)2 = (1/2)[OH-] = 2.64 * 10-6 M